OpenStudy (anonymous):

Simplify square root of 162 end root minus square root of 160 end root plus 5 square root of 32 end root minus square root of 90 negative 4 square root of 56 22 square root of 8 29 square root of 2 end root minus 7 square root of 10 29 square root of 4 end root minus 7 square root of 20

OpenStudy (anonymous):

\[\sqrt{162} - \sqrt{160} + 5\sqrt{32} - \sqrt{90}\]

OpenStudy (anonymous):

@lacrosseplayer22

OpenStudy (anonymous):

 so those are the simplified version of the roots. ill explain how I got those through the messaging but after that since only common roots can be added and subtractes we can do this: 9root2+20root2=29root2 -4root10-3root10=-7root10 so the final answer should be 29root2-7root2 which is the same thing as -7root2+29root2